The Ksp of Pb(IO3)2 is given as 20 * 10-13

1) Find the solublity of Pb(IO3)2 in water.

2) Find the solubility of Pb(IO3)2 in 1.0 * 10-4 M Pb(NO3)2.
Answer #1

1)

At equilibrium:

Pb(IO3)2 <----> Pb2+ + 2 IO3-

   s 2s

Ksp = [Pb2+][IO3-]^2

2*10^-13=(s)*(2s)^2

2*10^-13= 4(s)^3

s = 3.684*10^-5 M

Answer: 3.7*10^-5 M

2)

Pb(IO3)2 here is Strong electrolyte

It will dissociate completely to give [IO3-] = 1.0*10^-4 M

At equilibrium:

Pb(IO3)2 <----> Pb2+ + 2 IO3-

   s 1*10^-4 + 2s

Ksp = [Pb2+][IO3-]^2

2*10^-13=(s)*(1*10^-4+ 2s)^2

Since Ksp is small, s can be ignored as compared to 1*10^-4

Above expression thus becomes:

2*10^-13=(s)*(1*10^-4)^2

2*10^-13= (s) * 10^-8

s = 2.0*10^-5 M

Answer: 2.0*10^-5 M